# MuTox: Universal MUltilingual Audio-based TOXicity Dataset and Zero-shot Detector

### v1

@@ Abstract
Research in toxicity detection in natural language processing for the speech modality (audio-based) is quite limited, particularly for languages other than English. To address these limitations and lay the groundwork for truly multilingual audio-based toxicity detection, we introduce MuTox, the first highly multilingual audio-based dataset with toxicity labels. The dataset comprises 20,000 audio utterances for English and Spanish, and 4,000 for the other 19 languages. To demonstrate the quality of this dataset, we trained the MuTox audio-based toxicity classifier, which enables zero-shot toxicity detection across a wide range of languages. This classifier outperforms existing text-based trainable classifiers by more than 1% AUC, while expanding the language coverage more than tenfold. When compared to a wordlist-based classifier that covers a similar number of languages, MuTox improves precision and recall by approximately 2.5 times. This significant improvement underscores the potential of MuTox in advancing the field of audio-based toxicity detection.
@@ 1 Introduction
In this paper, we go far beyond existing research in audio-based toxicity detection by providing the first highly multilingual audio-based toxicity annotated dataset (MuTox dataset, 21 languages) together with the first text-less massive multilingual metric (MuTox classifier, 100+ languages). Note that multilinguality for audio-based toxicity detection becomes even more crucial for the task of added toxicity in the context of multimodal and multilingual translation, where the case of adding or deleting toxicity may be considered as a critical error Seamless Communication et al. (2023a).
In particular, the main contributions of this paper are: providing guidelines for audio-based toxicity annotation (section 3); releasing the first highly multilingual audio-based toxicity dataset and benchmark with human annotations for 21 languages (section 4); analyzing the performance of text-based classifiers when applied to audio-based toxicity detection (section 6); proposing MuTox, a massively multilingual audio-based toxicity classifier (section 5).
Our results can be summarized as follows:
  • When compared to the strongest performing systems, which are composed of speech recognition plus trainable text-based toxicity detectors, MuTox slightly surpasses (+1% on average), while offering more than 10 times the language coverage.
  • When compared to systems with the highest coverage, which are composed of speech recognition plus wordlists toxicity detectors, MuTox improves precision and recall by a factor of approximately 2.5.
@@ 2 Background
etox is an open-source [^2] https://github.com/facebookresearch/seamless_communnication/toxicity wordlist-based classifier covering 200 languages. While this metric has several advantages - namely, it covers a massive number of languages and is highly transparent - it has other limitations such as only detecting lexical toxicity and not distinguishing polysemous words which may be toxic in some contexts and not others.
detoxify is a text-based toxicity classifier trained mainly in JigSaw Kaggle (2018) which is freely available in 8 languages (English, French, German, Italian, Portuguese, Russian, Spanish, Turkish).
@@
We are not aware of any specific multilingual speech datasets which contain toxicity or toxicity labels. However, we want to pre-select data for annotation to maximize chances of the annotators confirming toxicity. Relatedly, Detoxy Ghosh et al. (2021) estimated the amount of toxicity for several English spoken datasets using text-based classifiers. In this work, we follow a similar approach, by exploring two highly multilingual spoken datasets: CommonVoice Ardila et al. (2020) and SeamlessAlign Seamless Communication et al. (2023b).
@@ 3 Annotation Guidelines
To help annotators reach a faster determination as to which words or phrases could qualify as toxic in this exercise, we also pointed them to the publicly available Toxicity-200333https://tinyurl.com/NLLB200TWL word list repository.
@@ 4 MuTox DataSet Description
MuTox is composed of 21 languages. Two languages, English and Spanish, contain a larger amount of annotated data (20k utterances/sentences each language), while the rest of languages [^4] Bengali, Dutch, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Mandarin Chinese, Arabic, Portuguese, Russian, Swahili, Tagalog, Thai, Turkish, Urdu, Vietnamese, contain a smaller amount of annotated data (4k utterances/sentences each language).
@@ 4.1 Annotation for English and Spanish
Annotation results are reported in Table 1. The percentages of confirmed toxic samples are 16 % and 19% for English and Spanish, respectively. Figure 1 reports the types of toxicity for each language using annotated categories.The proportion of toxicity increases with the toxicity quantile (see Figure 5 in appendix A) . The correlation coefficient between the toxicity in the annotation with detoxify threshold is 0.6.
@@
4.2 Annotation for 19 additional languages
@@
Based on annotated data in English and Spanish (section 4.1), we devise new criteria to select 4k sentences across 19 extra languages which we consider high-priority (HP) for the context of this work and related previous projects Seamless Communication et al. (2023a). The purpose of this extension is to form a highly multilingual benchmark for audio-based toxicity classification.
Languages. Bengali, Dutch, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Mandarin Chinese, Modern Standard Arabic, Portuguese, Russian, Swahili, Tagalog, Thai, Turkish, Urdu, Vietnamese.
@@
We include all the samples positive per etox and detoxify scores higher than 0.8. This results in a small number of samples, ranging from 502 (deu) to less than 10 (hin, tgl, cmn) Then, we include etox detections (ignoring the detoxify threshold) but allow only for a maximum of 200 sentences for each etox token and add a maximum of 1,000 etox sentences. We set these maxima so as to ensure toxicity diversity in the selection. Finally, we include samples detected with detoxify: we use detoxify in Italian, Portuguese, Turkish, Russian and French (which are supported by the tool). For the rest of non-supported languages, we resort to English detoxify in the parallel text. [This method is supported as follows: on the corpus from section 4.1, we have a total toxicity of 0.25 in Spanish when using detoxify in English with a threshold of 0.8].
@@
Figure 2 shows the results of this selection, grouped by the methodology used (1 - intersection, 2 - etox, 3 - detoxify). Table 2 shows the dataset splits. The final dataset of 19 languages includes 120 hours.
@@ 5 MuTox audio-based toxicity Classifier
We feed our toxicity classifier, MuTox, with both audio-based and text-based toxicity-labeled data. The audio-based toxicity classifier depicted on 3 follows a simple architecture consisting of an encoder, turning input text or audio speech into a fixed-size representation vector, and a binary classifier composed of three simple feed-forward layers.
@@
We use multimodal and multilingual SONAR encoders from Duquenne et al. (2023). This choice is motivated by the broad language coverage at the time of this study (English, Spanish, and HP languages) and the zero-shot capabilities of SONAR: once trained on a set of languages, the classifier head can be used on top of any compatible SONAR encoder handling another language. For the classifier, we use variable input sizes for the 3 feedforward layers (1024, 512, and 128). We train with a Binary Cross Entropy loss with logits and Adam optimizer with an initial learning rate of 0.001. In order to compare zero-shot (ZS) vs supervised performance, we train a first classifier with English and Spanish training data only and test on HP languages. We then train a second classifier on all training data available and report results on the same HP languages. Results are listed in Table 2. Throughout our experiments, we use Whisper-Large-v2 to generate transcriptions.
@@ 6 Experiments and Results
We evaluate MuTox (text and speech) on Devtest and Test from Table 2. We compare the performance against ASR-detoxify (hereinafter, detoxify for simplicity) using available tool1 and ASR-etox (hereinafter, etox) using available tool2. ASR of the speech is done with Whisper-Large-v2. We report Area Under the Curve (AUC), Precision and Recall across toxicity detection.
Correlation accross classifiers
We computed Pearson correlation across classifiers see results in Table 3. To compute the correlations we computed logarithm to MuTox outputs. We observe that correlation is higher between MuTox and detoxify than MuTox and etox or even detoxify and etox. In future work, we would like to do a manual analysis of the differences across classifiers.
ASR-MuTox MuTox detoxify etox
ASR-MuTox 1 0.75 0.51 0.13
MuTox 0.75 1 0.49 0.07
detoxify 0.51 0.49 1 0.26
etox 0.13 0.07 0,26 1
Table 3: Pearson correlation accross toxicity classifiers.
@@
Table 4 compares MuTox and Detoxify (for available languages) in terms of AUC. When comparing the 8 languages covered by detoxify, MuTox trained on all languages shows a slight improvement over Detoxify of 1%. MuTox however, scales to 10 times more languages than detoxify.
@@
Supervised MuTox improves zero-shot MuTox by more than 7% on average. When comparing MuTox and ASR-MuTox averaging over 21 languages, results are comparable for supervised training. While it is unclear why MuTox vs. ASR-MuTox show different results depending on the language, we could hypothesize that the imbalances in the complexities of pronunciation/writing in various languages lead to variations in transcription quality. ASR-MuTox recall (at fixed precision) is higher than when using MuTox directly on audio speech (by more than 8%). Note that zero-shot MuTox outperforms on average zero-shot ASR-MuTox. On test, we get average 0.36 recall at fixed precision compared to 0.27, and 0.76 AUC compared to 0.73.
[…]
@@
lang etox MuTox-ZS ASR-MuTox-ZS MuTox ASR-MuTox
@@
eng 0.39 0.30 0.40 0.31 - - - - 0.61 0.18 0.75 0.49
spa 0.40 0.33 0.41 0.33 - - - - 0.64 0.19 0.73 0.44
arb 0.09 0.02 0.18 0.03 0.16 0.26 0.18 0.15 0.57 0.58 0.40 0.64
ben 0.02 0.02 0.01 0.02 0.35 0.11 0.10 0.03 0.55 0.40 0.15 0.37
cmn 0.00 0.00 0.00 0.00 0.14 0.37 0.27 0.17 0.36 0.32 0.23 0.13
deu 0.42 0.36 0.43 0.37 0.97 0.79 0.95 0.80 0.91 0.91 0.88 0.90
fra 0.11 0.42 0.10 0.40 0.26 0.31 0.08 0.14 0.26 0.62 0.57 0.58
hin 0.18 0.02 0.09 0.01 0.56 0.36 0.54 0.33 0.63 0.77 0.71 0.86
ind 0.10 0.22 0.12 0.30 0.18 0.19 0.07 0.18 0.44 0.43 0.22 0.39
ita 0.20 0.38 0.18 0.32 0.10 - 0.28 0.18 0.10 0.12 0.35 0.33
kor 0.02 0.13 0.00 0.02 - - - - - - - -
nld 0.08 0.15 0.06 0.11 0.77 0.73 0.21 0.21 0.79 0.88 0.62 0.56
por 0.29 0.43 0.28 0.45 0.68 0.69 0.75 0.77 0.74 0.67 0.88 0.75
rus 0.17 0.44 0.18 0.46 0.54 0.42 0.47 0.52 0.78 0.43 0.77 0.69
swh 0.08 0.18 0.07 0.17 0.04 0.02 - 0.06 0.02 - - -
tgl 0.27 0.09 0.14 0.04 0.19 0.05 0.05 0.06 0.35 0.34 0.30 -
tha 0.00 0.00 0.00 0.00 - - 0.14 0.07 - - 0.14 -
tur 0.06 0.08 0.10 0.13 0.13 - 0.15 0.12 0.34 0.37 0.43 0.47
vie 0.12 0.16 0.10 0.16 0.54 0.69 0.46 0.61 0.74 0.77 0.71 0.60
avg 0.16 0.20 0.15 0.19 - - - - 0.46 0.50 0.49 0.55
Table 5: Toxicity detection precision and recall results. MuTox recall at the precision of $max(\textsc{etox}-precision,0.3)$ vs. etox. The best results are bolded..
@@
Figure 4 reports the model performances by toxicity categories. We note that detoxify and MuTox have less variance (<0.001) across categories than etox (>0.01). etox performs well on Profanities (recall>0.8), since those can easily be identified with explicit wordlists. It struggles however with more complex/implicit types of toxicity such as Hate Speech or Physical Violence (recall shows to be below 0.2 for these).

### v2

@@ Abstract
Research in toxicity detection in natural language processing for the speech modality (audio-based) is quite limited, particularly for languages other than English. To address these limitations and lay the groundwork for truly multilingual audio-based toxicity detection, we introduce MuTox, the first highly multilingual audio-based dataset with toxicity labels which covers 14 different linguistic families. The dataset comprises 20,000 audio utterances for English and Spanish, and 4,000 for the other 28 languages. To demonstrate the quality of this dataset, we trained the MuTox audio-based toxicity classifier, which enables zero-shot toxicity detection across a wide range of languages. This classifier performs on par with existing text-based trainable classifiers, while expanding the language coverage more than tenfold. When compared to a wordlist-based classifier that covers a similar number of languages, MuTox improves F1-Score by an average of 100%. This significant improvement underscores the potential of MuTox in advancing the field of audio-based toxicity detection.
@@ 1 Introduction
In this paper, we go far beyond existing research in audio-based toxicity detection by providing the first highly multilingual audio-based toxicity annotated dataset (MuTox dataset, 30 languages, see Table 6 in appendix A) together with the first text-less massive multilingual metric (MuTox classifier, 100+ languages). Note that multilinguality for audio-based toxicity detection becomes even more crucial for the task of added toxicity in the context of multimodal and multilingual translation, where the case of adding or deleting toxicity may be considered as a critical error Seamless Communication et al. (2023a).
In particular, the main contributions of this paper are: providing guidelines for audio-based toxicity annotation (section 3); releasing the first highly multilingual audio-based toxicity dataset and benchmark with human annotations for 30 languages (section 4); analyzing the performance of text-based classifiers when applied to audio-based toxicity detection (section 6); proposing MuTox, a massively multilingual audio-based toxicity classifier (section 5). Our results can be summarized:
  • When compared to the strongest performing systems, which are composed of speech recognition plus trainable text-based toxicity detectors, MuTox performs on par, while offering more than 10 times the language coverage  [^2] We want to clarify that while MuTox dataset evaluates 30 languages, the MuTox classifier relies on SONAR Duquenne et al. (2023) embeddings. As of March 2024, SONAR encoders are available for 57 languages in speech and 200 in text. Thanks to this architecture, by design, the MuTox classifier provides zero-shot toxicity detection for all of them. This is more than 8 times language coverage in speech and more than 25 times in text..
  • When compared to systems with the highest coverage, which are composed of speech recognition plus wordlists toxicity detectors, MuTox improves F1-Score by an average of 100%.
@@ 2 Background
etox is an open-source [^3] https://github.com/facebookresearch/seamless_communication/tree/mutox/src/seamless_communication/cli/toxicity/mutox wordlist-based classifier covering 200 languages. While this metric has several advantages - namely, it covers a massive number of languages and is highly transparent - it has other limitations such as only detecting lexical toxicity and not distinguishing polysemous words which may be toxic in some contexts and not others.
detoxify 1 is a text-based toxicity classifier trained mainly in JigSaw Kaggle (2018) which is freely available in 7 languages (English, French, Italian, Portuguese, Russian, Spanish, Turkish).
@@
There are extremely few speech datasets with toxicity labels. Recently, ADIMA covers multilingual profanity detection audio dataset in 10 Indic languages Gupta et al. (2022). Relatedly, Detoxy Ghosh et al. (2021) estimated the amount of toxicity for several English spoken datasets using text-based classifiers. In this work, we follow a similar approach for pre-selecting data for annotation to maximize chances of the annotators confirming toxicity. We used two spoken datasets which cover diverse domains and are highly multilingual: CommonVoice Ardila et al. (2020) and SeamlessAlign Seamless Communication et al. (2023b).
@@ 3 Annotation Guidelines
To help annotators reach a faster determination as to which words or phrases could qualify as toxic in this exercise, we also pointed them to the publicly available Toxicity-200444https://tinyurl.com/NLLB200TWL word list repository.
@@ 4 MuTox DataSet Description
MuTox is composed of 30 languages. Two languages, English and Spanish, contain a larger amount of annotated data (20k utterances/sentences each language), while the rest of languages (see section 4.2), contain a smaller amount of annotated data (4k utterances/sentences each language).
@@ 4.1 Annotation for English and Spanish
Annotation results are reported in Table 1. The percentages of confirmed toxic samples are 16 % and 19% for English and Spanish, respectively. Figure 1 reports the types of toxicity for each language using annotated categories.The proportion of toxicity increases with the toxicity quantile (see Figure 5 in appendix B) . The correlation coefficient between the toxicity in the annotation with detoxify threshold is 0.6.
@@
4.2 Annotation for 28 additional languages
@@
Based on annotated data in English and Spanish (section 4.1), we devise new criteria to select 4k sentences across 28 extra languages which we consider high-priority (HP) for the context of this work and related previous projects Seamless Communication et al. (2023a). The purpose of this extension is to form a highly multilingual benchmark for audio-based toxicity classification.
Languages. Modern Standard Arabic, Bengali, Bulgarian, Catalan, Czech, Danish, Dutch, Estonian, Finish, French, German, Greek, Hebrew, Hindi, Hungarian, Indonesian, Italian, Mandarin Chinese, Persian, Polish, Portuguese, Russian, Slovak, Swahili, Tagalog, Turkish, Urdu, Vietnamese (see table 6).
@@
We include all the samples positive per etox and detoxify scores higher than 0.8 (i.e. intersection). This results in a small number of samples, ranging from more than 2k (pol) to 0 (cmn,jpn). Then, we include etox detections (ignoring the detoxify threshold) but allow only for a maximum of 200 sentences for each etox token and add a maximum of 1,000 etox sentences. We set these maxima so as to ensure toxicity diversity in the selection. Finally, we include samples detected with detoxify: we use detoxify in Italian, Portuguese, Turkish, Russian and French (which are supported by the tool). For the rest of non-supported languages, we resort to English detoxify in the parallel text. The effectiveness of this method is supported by the following evidence: when using detoxify in English with a threshold of 0.80 on the corpus from section 4.1, we observe a total toxicity of 0.25 in Spanish.
@@
Figure 2 shows the results of this selection, grouped by the methodology used (1 - intersection, 2 - etox, 3 - detoxify). Table 2 shows the dataset splits. The final dataset of 28 languages includes 83 hours.
@@ 5 MuTox audio-based toxicity Classifier
We feed our toxicity classifier, MuTox, with both audio-based and text-based toxicity-labeled data. The audio-based toxicity classifier depicted on Figure 3 follows a simple architecture consisting of an encoder, turning input text or audio speech into a fixed-size representation vector, and a binary classifier composed of three simple feed-forward layers.
@@
We use multimodal and multilingual SONAR encoders from Duquenne et al. (2023). This choice is motivated by the broad language coverage at the time of this study (English, Spanish, and HP languages) and the zero-shot capabilities of SONAR: once trained on a set of languages, the classifier head can be used on top of any compatible SONAR encoder handling another language. For the classifier, we use variable input sizes for the 3 feedforward layers (1024, 512, and 128). We train with a Binary Cross Entropy loss with logits and Adam optimizer with an initial learning rate of 0.001. In order to compare zero-shot (ZS) vs supervised performance, we train a first classifier with English and Spanish training data only and test on HP languages. We then train a second classifier on all training data available and report results on the same HP languages. The number of parameters of our models is around 600k. Results are listed in Table 4. Throughout our experiments, we use Whisper-Large-v2 to generate transcriptions.
@@ 6 Experiments and Results
We evaluate MuTox (text and speech) on Devtest and Test from Table 2. We compare the performance against ASR-detoxify (hereinafter, detoxify for simplicity) using the available tool1 and ASR-etox (hereinafter, etox) using the available tool3. Speech recognition is done with Whisper-Large-v2. We report Area Under the Curve (AUC), Precision, Recall and F1-Score across toxicity detection.
Correlation across classifiers
We computed Pearson correlation across classifiers (see results in Table 3). We observe that correlation is higher between MuTox and detoxify than MuTox and etox or even detoxify and etox. In future work, we would like to do a manual analysis of the differences across classifiers.
etox detoxify MuTox ASR-MuTox
etox 1 0.26 0.06 0.12
detoxify 0.26 1 0.46 0.49
MuTox 0.06 0.46 1 0.79
ASR-MuTox 0.12 0.49 0.79 1
Table 3: Pearson correlation across toxicity classifiers.
@@
Table 4 compares MuTox and Detoxify (for available languages) in terms of AUC. When comparing the 7 languages covered by detoxify, MuTox trained on all languages shows on par results with detoxify. MuTox however, scales to 10 times more languages than detoxify.
@@
Supervised MuTox slightly improves zero-shot MuTox by 2% on average. When comparing MuTox and ASR-MuTox averaging over 30 languages, results are almost comparable for zero-shot, but ASR-Mutox is better than MuTox in supervised setting. While it is unclear why MuTox vs. ASR-MuTox show different results depending on the language, we could hypothesize that the imbalances in the complexities of pronunciation/writing in various languages lead to variations in transcription quality.
lang Devtest (D) Test (T) MuToxZS ASR-MuToxZS MuTox ASR-MuTox detoxify
@@
lang etox MuToxZS ASR-MuToxZS MuTox ASR-MuTox
@@
eng 0.39 0.30 0.40 0.31 - - - - 0.18 0.23 0.58 0.67
spa 0.40 0.33 0.41 0.33 - - - - 0.19 0.32 0.73 0.74
arb 0.09 0.02 0.21 0.04 0.16 0.26 0.17 0.15 0.60 0.54 0.44 0.53
ben 0.02 0.05 0.00 0.00 0.35 0.11 0.10 0.03 0.30 0.45 0.45 0.50
bul 0.19 0.29 0.17 0.21 0.22 0.32 0.53 0.65 0.71 0.69 0.74 0.65
cat 0.09 0.36 0.08 0.30 0.18 0.07 0.18 0.19 0.59 0.05 0.77 0.16
ces 0.12 0.44 0.18 0.73 0.19 0.45 0.30 0.59 0.21 0.41 0.30 0.61
cmn 0.00 0.00 0.00 0.00 0.14 0.37 0.27 0.17 0.43 0.35 0.41 0.16
dan 0.26 0.77 0.23 0.68 0.92 0.91 0.93 0.90 0.89 0.92 0.96 0.88
deu 0.42 0.36 0.43 0.36 0.81 0.88 0.95 0.99 0.86 0.90 0.98 0.99
ell 0.11 0.39 0.12 0.41 0.02 0.16 0.03 0.36 0.19 0.18 0.44 0.31
est 0.15 0.49 0.15 0.51 0.57 0.54 0.71 0.63 0.86 0.84 0.80 0.76
fas 0.07 0.75 0.07 0.68 0.15 0.23 0.18 0.33 0.33 0.14 0.50 0.36
fin 0.17 0.90 0.17 0.89 0.91 0.86 0.94 0.94 0.87 0.88 0.94 0.94
fra 0.11 0.42 0.10 0.39 0.26 0.31 0.08 0.14 0.40 0.53 0.48 0.50
heb 0.05 0.63 0.04 0.53 0.10 0.05 0.07 0.10 0.13 0.05 0.03 0.70
hin 0.18 0.02 0.06 0.01 0.56 0.36 0.54 0.33 0.63 0.72 0.66 0.79
hun 0.20 0.73 0.22 0.79 0.34 0.51 0.57 0.57 0.55 0.49 0.56 0.64
ind 0.10 0.22 0.11 0.28 0.18 0.19 0.07 0.17 0.15 0.41 0.32 0.40
ita 0.20 0.38 0.16 0.27 0.10 0.05 0.28 0.18 0.16 0.21 0.31 0.35
nld 0.08 0.15 0.06 0.11 0.77 0.73 0.21 0.21 0.78 0.84 0.61 0.52
pol 0.12 0.89 0.12 0.88 0.86 0.60 0.83 0.73 0.67 0.73 0.86 0.79
por 0.28 0.43 0.30 0.50 0.68 0.69 0.75 0.77 0.74 0.59 0.66 0.63
rus 0.17 0.44 0.19 0.47 0.54 0.42 0.47 0.52 0.65 0.27 0.69 0.63
slk 0.14 0.73 0.09 0.48 0.49 0.41 0.36 0.52 0.47 0.56 0.49 0.66
swh 0.08 0.18 0.07 0.15 0.04 0.02 0.64 0.06 0.02 0.61 0.60 0.46
tgl 0.27 0.09 0.12 0.03 0.19 0.05 0.05 0.06 0.35 0.23 0.07 0.09
tur 0.06 0.08 0.12 0.16 0.13 0.93 0.15 0.12 0.34 0.28 0.28 0.47
urd 0.00 0.00 0.00 0.00 0.87 0.87 0.81 0.85 0.88 0.88 0.83 0.92
vie 0.12 0.16 0.10 0.15 0.54 0.69 0.46 0.61 0.68 0.80 0.73 0.68
avg 0.17 0.36 0.17 0.35 0.40 0.43 0.42 0.42 0.47 0.48 0.57 0.58
Table 5: Toxicity detection precision and recall results. MuTox recall at the precision of $max(\textsc{etox},0.3)$ vs. etox. The best results are bolded.
@@
Figure 4 reports the model performances by toxicity categories. We note that detoxify and MuTox have less variance (<0.001) across categories than etox (>0.01). etox performs well on Profanities (recall>0.8), since those can easily be identified with explicit wordlists. It struggles however with more complex/implicit types of toxicity such as Hate Speech or Physical Violence.